# Enhancing Automated Loop Invariant Generation for Complex Programs with Large Language Models

### v1

@@ Abstract
Automated program verification has always been an important component of building trustworthy software. While the analysis of real-world programs remains a theoretical challenge, the automation of loop invariant analysis has effectively resolved the problem. However, real-world programs that often mix complex data structures and control flows pose challenges to traditional loop invariant generation tools. To enhance the applicability of invariant generation techniques, we proposed ACInv, an Automated Complex program loop Invariant generation tool, which combines static analysis with Large Language Models (LLMs) to generate the proper loop invariants. We utilize static analysis to extract the necessary information for each loop and embed it into prompts for the LLM to generate invariants for each loop. Subsequently, we employ an LLM-based evaluator to assess the generated invariants, refining them by either strengthening, weakening, or rejecting them based on their correctness, ultimately obtaining enhanced invariants. We conducted experiments on ACInv, which showed that ACInv outperformed previous tools on data sets with data structures, and maintained similar performance to the state-of-the-art tool AutoSpec on numerical programs without data structures. For the total data set, ACInv can solve 21% more examples than AutoSpec and can generate reference data structure templates.
Enhancing Automated Loop Invariant Generation for Complex Programs with Large Language Models
Ruibang Liu, Guoqiang Li, Minyu Chen, Ling-I Wu, Jingyu Ke Shanghai Jiao Tong University, Shanghai, 200240, China {628628,li.g, minkow,wulianyi1998, Windocotber}@sjtu.edu.cn
@@ 1 Introduction
Program verification has progressively emerged as a fundamental component in the development of reliable software within the industry, thereby necessitating advanced tools for automated program verification. Nonetheless, the verification of program properties in the presence of loops is inherently undecidable. Within the framework of Hoare logic Hoare (1969), loop invariants, which are usually assertions that are always true during the entry and execution of the loop, serve as the abstractions of loop properties. However, in the practical implementation of program verification tools such as CPAchecker Beyer and Keremoglu (2011), CBMC Kroening and Tautschnig (2014), SMACK Rakamarić and Emmi (2014) and Frama-C Kirchner et al. (2015), the derivation of loop invariants often necessitates manual intervention by domain experts, which poses a substantial impediment to the full automation of program verification.
@@
Loop invariants are usually defined inductively, which means that they must meet the following three conditions: (1) Initialization: It is true before the first iteration of the loop; (2) Maintenance: If the invariant holds true at the beginning of an iteration, it must also hold true at the end of that iteration. Abstract interpretation-based methods Rodríguez-Carbonell and Kapur (2004); Rodríguez-Carbonell and Kapur (2007); de Oliveira et al. (2016) was initially adopted by researchers to automated loop invariants generation. These approaches typically employ static analysis techniques, such as value analysis and interval analysis, to trace the values manipulated by the program and subsequently synthesize these values to construct loop invariants. However, these methods are limited to analyzing programs with simple numerical types and fail to adequately address intricate relationships between variables. Template-based methods Sankaranarayanan et al. (2004); Yao et al. (2023a); Liu and Li (2024); Colón et al. (2003); Liu et al. (2022) typically involve predefining invariants using SMT formulas, proposing parameterized invariant templates via uninterpreted functions or inequalities, and employing heuristic strategies to determine the parameters. While this approach can reveal more complex relationships between variables, it remains largely confined to linear numerical inequalities and imposes significant computational costs due to the introduction of SMT constraint solving. The widespread adoption of machine learning has led many researchers to propose learning-based methods Ryan et al. (2020); Garg et al. (2014); Si et al. (2018); Yu et al. (2023), which infer suitable invariant templates through specific deductions and learn the corresponding parameters using various learning techniques to generate invariants. However, their effectiveness remains limited by the predefined templates and the quality of the training data. Although there is a substantial body of prior work on the automatic generation of loop invariants, these methods often face challenges when applied to the verification of real-world programs, primarily due to their generality and associated performance overhead.
@@
In recent years, Large Language Models (LLMs), such as ChatGPT, have emerged as powerful tools for automating various generative tasks. Several studies Li et al. (2024); Su et al. (2024) have demonstrated their strong performance in program-related reasoning tasks. Notably, recent works Chakraborty et al. (2023); Pei et al. (2023); Wen et al. (2024); Liu et al. (2024); Kamath et al. (2023) have highlighted the potential of LLMs in generating loop invariants. LLMs excel at recognizing complex structures, such as loops, and are capable of automating invariant generation across extensive codebases, significantly reducing the need for manual effort. On the other hand, as noted in previous work Liu et al. (2023); Wei et al. (2022); Yao et al. (2023b); Yang et al. (2024), LLMs face certain challenges in invariant generation. First, LLMs may produce incorrect invariants, and these errors can accumulate when the models are invoked repeatedly. Second, without sufficient information or carefully crafted prompts, even for correct invariants, LLMs tend to generate weaker invariant constraints, which may prevent the verification of the program’s target properties.
To address these challenges, we propose ACInv, an Automated Complex program loop Invariant generation tool leveraging LLMs in conjunction with static analysis. Specifically, we argue that one of the complexities of real-world programs arises from the variety of data types and user-defined data structures, as well as the complexity of loop constructs. Consequently, ACInv must first leverage static analysis techniques to extract information on both the data structures and control flow of the program, and the information obtained can enhance the LLM’s understanding of the program. Furthermore, it is crucial to provide the LLM with opportunities to assess and revise the generated invariants, which not only corrects erroneous invariants but also attempts to strengthen the valid ones. Therefore, ACInv, after leveraging static analysis to extract essential loop information, generates predicates that abstract the properties of data structures used in the program and subsequently generate the corresponding invariants. Once the invariants are generated, ACInv employs an LLM-based evaluator to assess their correctness. Based on this evaluation, the LLM further optimizes the invariants. For correct invariants, ACInv attempts to strengthen them, refining their scope. For incorrect invariants, ACInv weakens them to encompass the correct invariant space. If even after weakening, the invariant cannot cover the correct space, the invariant is rejected.
@@
We contend that ACInv offers several distinct advantages over existing tools. First, it is capable of analyzing a wide range of complex, real-world code, with its sophisticated template processing for data structures and effective handling of nested loops enabling it to tackle intricately structured programs. Second, ACInv delivers a fully automated, end-to-end solution for loop invariant generation, significantly reducing the need for manual intervention throughout the process. Lastly, while maintaining a competitive level of accuracy, ACInv minimizes time overhead and optimizes the quality of the generated invariants, thereby enhancing their precision and robustness in practical applications. To evaluate ACInv, we conducted a series of experiments. We conducted experiments on ACInv on datasets with and without data structures. The experiments showed that ACInv performed better than all current tools on datasets with data structures. On datasets without data structures, ACInv also performed close to the performance of the state-of-art tool AutoSpec Wen et al. (2024). Overall, ACInv can solve 21.08% more examples than AutoSpec.
@@
  • We constructed a dataset that is more representative of real-world programs for loop invariants
  • We generated some predicate templates for some commonly used data structures and other user-defined data structures.
  • We developed a more realistic, real-world program-oriented automatic loop invariant generation framework and implemented it as a tool ACInv.
@@ 2.1 Loop Invariant
To provide a formal characterization of the invariant, we first abstract a loop of programs as follows:
$Pre(X)\ while(B)\{S\}\ Post(X)$
where $X$ is the set of variables occurring in the program, $Pre(X)$ and $Post(X)$ denote the pre-conditions of the loop and post-conditions of the loop, $B$ represents the loop condition, and $S$ is the loop body.
Traditional symbolic methods for invariant generation typically involve direct code analysis, mainly including abstraction interpretation and constraint solving. Abstract interpretation Cousot and Cousot (1977) approximates a program’s behavior by mapping its concrete semantics to a simplified abstract domain, often using fixed-point calculations to abstract program states in the presence of loops, such as Enric et al.Rodríguez-Carbonell and Kapur (2004) and Steven et al. de Oliveira et al. (2016). Constraint-based methods, on the other hand, rely on pre-defined templates and employ SMT solvers to filter out appropriate invariant expressions. Farkas’ lemma Sankaranarayanan et al. (2004); Yao et al. (2023a); Liu and Li (2024); Colón et al. (2003); Liu et al. (2022) is frequently used to extract linear relationships from loops, while other constraint templates, such as those based on Craig’s interpolant McMillan (2003); Dillig et al. (2013a, b), are also employed.
Learning-based methods often build upon template-based approaches by learning the parameters within these templates through various machine-learning techniques and ultimately invoking an SMT solver to filter suitable invariants. C2I Sharma and Aiken (2014) employs a randomized search to discover candidate invariants, followed by a checker to validate them. ICE Garg et al. (2014) did learning for synthesizing invariants, leveraging examples, counterexamples, and implications. Then they combined them with decision trees Garg et al. (2016). CODE2INV Si et al. (2018, 2020) utilizes reinforcement learning and graph neural networks to learn program structures. LIPuS Yu et al. (2023) also adopts reinforcement learning and uses pruning to reduce the search space. CLN2INV Ryan et al. (2020) introduces Continuous Logic Networks (CLN) to automatically learn loop invariants directly from program execution traces, while Gated-CLN (G-CLN) Yao et al. (2020) extends this approach to allow the model to robustly learn general invariants across a large number of terms.
@@ 2.2 Large Language Models (LLMs)
In recent years, LLMs have emerged as powerful tools for understanding and generating human-like language. The development of large-scale pre-trained models like GPT (Generative Pre-trained Transformer) Brown et al. (2020) has led to substantial improvements across a variety of NLP tasks including code-related task with prompting skills like Chain-of-Thought (CoT) prompting Wei et al. (2022) and few-shot learning Brown et al. (2020).
Several works have explored the intersection of LLMs and loop invariant generation with promising outcomes. Autospec Wen et al. (2024) employs hierarchical specification generation to produce ACSL-style annotations, which include loop invariants, and has achieved good performance in the processing of numerical programs. LIG-SE Liu et al. (2024) fine-tuned LLMs and combining it with a checker gives the large model the ability to handle separation logic. It can also handle data structures and abstract predicate properties to a certain extent. The iRank tool Chakraborty et al. (2023) introduces a re-ranking approach that evaluates loop invariants produced by LLMs. Unlike other methods, iRank does not focus on generating loop invariants but instead addresses the challenge of verifying the correctness of LLM-generated invariants. To achieve this, the authors employed two distinct LLM-based embedding models to learn the criteria for determining invariant correctness. Kexin et al.Pei et al. (2023) aimed to fine-tune a pre-trained LLM for abstract reasoning over program executions, allowing for the generation of loop invariants. Their approach relies on a dynamic analyzer to trace program executions, which are then used to derive invariants. However, their evaluation is limited to a comparison against invariants generated by the dynamic analyzer, which poses inherent constraints. LoopyKamath et al. (2023), on the other hand, continuously queries LLMs to generate invariants and employs a verification tool to ensure correctness. While they enhance the generated invariants by cyclically applying the Houdini algorithm and making repairs, the repeated invocation of the verifier introduces significant performance overhead.
@@ 3.1 Overview
In this section, we will first provide an overview of the entire workflow, followed by a detailed explanation of each component. In Figure 1, we present an overview diagram illustrating our proposed approach. To address the aforementioned challenge, we introduce ACInv, an LLM-based tool for automatic loop invariant generation. ACInv’s process consists of three main phases: (1) Extraction Phase: ACInv initially extracts data structure and loop information from the target program. This involves performing basic control flow and value analyses, classifying loops, and gathering relevant variable information for each loop. (2) Generation Phase: During this phase, ACInv embeds the extracted data structure information into data structure prompts, querying LLMs to derive predicates that abstract key properties of the data structure. These predicates, along with loop-specific information, are then embedded into invariant prompts to generate loop invariants. If the current loop is not the first in the program, information from preceding loops, along with their generated invariants, will also be incorporated into the prompts. (3) Augmentation Phase: This section mainly consists of two components: the evaluator, and the optimizer, which are all composed of LLMs. The evaluator assesses the correctness of each generated invariant. For correct invariants, the optimizer attempts to strengthen them to enhance precision. For incorrect invariants, the optimizer weakens them to align with the correct program state. If an invariant cannot be refined to the correct level after $k$ iterations, it is discarded. In the remainder of this section, we will provide a detailed explanation of the entire methodology and its related intricacies.
3.2 Prompts Design

### v3

@@ Abstract
Automated program verification has always been an important component of building trustworthy software. While the analysis of loops remains a theoretical challenge, the automation of loop invariant analysis has effectively resolved the problem. However, existing invariant generation tools are predominantly effective for programs with purely numerical or purely pointer-based structures. Real-world programs often mix complex data structures and control flows. These structures can include arrays, pointers, and recursive definitions, while control flows may involve multiple nested or concurrent loops. Traditional methods generally only generate invariants for simple numerical programs or specific segments, lacking broad applicability. In order to automatically generate loop invariants for real-world programs, we proposed ACInv, an Automated Complex program loop Invariant generation tool, which combines static analysis with prompting with Large Language Models (LLM) to generate the proper loop invariants. We employ static analysis to systematically decompose the program’s data structures and loops. This involves layer-by-layer transmission of structural information about variables, numerical data, and the complete loop structure to the LLM, enabling the generation of corresponding invariants. In comparison to prior work on AutoSpec, we delve deeper into the variable information within each loop. We conducted experiments on ACInv, which showed that ACInv outperformed previous tools on data sets with data structures and maintained similar performance to the state-of-the-art tool AutoSpec on numerical programs without data structures. For the total data set, ACInv can solve 21% more examples than AutoSpec, and can generate reference data structure templates.
keywords:
loop invariant generation, static analysis, large language models
journal: Science of Computer Programming
[…]
[sjtu]organization=Shanghai Jiao Tong University, postcode=200240, state=Shanghai, country=China
@@ 1 Introduction
Program verification has progressively emerged as a fundamental component in the development of reliable software within the industry, thereby necessitating advanced tools for automated program verification. Nonetheless, the verification of program properties in the presence of loops is inherently undecidable. Within the framework of Hoare logic [1], loop invariants, which are usually assertions that are always true during the entry and execution of the loop, serve as critical abstractions of loop properties. These invariants are indispensable in enabling verification tools to rigorously establish the correctness of programs. However, in the practical implementation of program verification tools such as CPAchecker [2], CBMC [3], SMACK [4] and Frama-C [5], the derivation of loop invariants often necessitates manual intervention by domain experts, which poses a substantial impediment to the full automation of program verification.
Loop invariants are usually defined inductively, which means that they must meet the following two conditions: (1) Initialization: It is true before the first iteration of the loop; (2) Maintenance: If the invariant holds true at the beginning of an iteration, it must also hold true at the end of that iteration. From a theoretical perspective, for any real-world program containing loops, irrespective of the programming language or the complexity of the program’s code structure or data structure, we aim to automatically generate correct loop invariants. Moreover, these loop invariants should be strong enough to support the verification of relevant properties under various conditions.
@@
Although there is a substantial body of prior work on the automatic generation of loop invariants, these methods often face challenges when applied to the verification of real-world programs, primarily due to their generality and associated performance overhead. Abstract interpretation baesd methods [6, 7, 8] was initially adopted by researchers. These approaches typically employ static analysis techniques, such as value analysis and interval analysis, to trace the values manipulated by the program and subsequently synthesize these values to construct loop invariants. However, these methods are limited to analyzing programs with simple numerical types and are less effective in handling complex loops, advanced data types and structures, and fail to adequately address intricate relationships between variables. Template-based methods [9, 10, 11, 12, 13] typically involve predefining invariants using SMT formulas, proposing parameterized invariant templates via uninterpreted functions or inequalities, and employing heuristic strategies to determine the parameters. While this approach can reveal more complex relationships between variables, it remains largely confined to linear numerical inequalities and imposes significant computational costs due to the introduction of SMT constraint solving. The widespread adoption of machine learning has led many researchers to propose learning-based methods [14, 15, 16, 17], which infer suitable invariant templates through specific deductions and learn the corresponding parameters using various learning techniques to generate invariants. However, their effectiveness remains limited by the predefined templates and the quality of the training data.
@@
In recent years, Large Language Models (LLMs), such as ChatGPT, have emerged as powerful tools for automating various generative tasks. Several studies [18, 19] have demonstrated their strong performance in program-related reasoning tasks. Notably, recent works [20, 21, 22, 23, 24] have highlighted the potential of LLMs in generating loop invariants. Trained on vast datasets encompassing both code and natural language, LLMs possess a sophisticated understanding of code patterns and semantic relationships. They excel at recognizing complex structures, such as loops, and are capable of automating invariant generation across extensive codebases, significantly reducing the need for manual effort.
Despite the potential of LLMs in invariant generation, several challenges remain. First, as highlighted by prior studies [25, 26, 27, 28], the inferential capabilities of LLMs require further exploration, particularly through carefully crafted prompts. Second, LLMs may produce incorrect invariants, and these errors can compound when the models are repeatedly invoked. To mitigate this issue, researchers have integrated SMT solvers or verification tools into the workflow to filter out erroneous invariants. However, this solution introduces significant performance overhead, and its effectiveness hinges on the strength of the verification tool. Moreover, certain correct yet complex loop invariants may be rejected by the tool, necessitating additional manual intervention, which conflicts with the goal of an end-to-end approach. Lastly, even when error-free, the invariants generated by LLMs may lack sufficient strength to capture precise properties, thereby failing to assist in verifying the target property.
To address these challenges, we propose ACInv, an Automated Complex program loop Invariant generation tool leveraging Large Language Models (LLMs) in conjunction with static analysis. In order for ACInv to be able to abstract data structures and reason about invariants, we need some necessary program information. The tool begins by extracting detailed information about the program’s data structures and variable values within loops using static analysis techniques. ACInv then queries LLMs to generate appropriate predicate templates, which abstract potential data structure properties in the program. After this step, it inserts the predicates generated by LLMs as comments in the program, further analyzing variable information within loops and annotating those variables that undergo value changes in each iteration. For nested loops, the tool recursively queries LLMs, starting with the innermost loop.
In order to enable ACInv to determine the correctness of the invariants it generates and further enhance the invariants it generates, we check and enhance each invariant generated by ACInv. This step is completed by the evaluator and optimizer built by LLM. The evaluator first evaluates the correctness of the invariant and then passes the evaluation results and the corresponding invariant to the optimizer for processing. For valid invariants, the optimizer seeks to refine them, while incorrect invariants are either weakened to align with the correct program states or discarded altogether.
@@
We contend that ACInv offers several distinct advantages over existing tools. First, it is capable of analyzing a wide range of complex, real-world code, with its sophisticated template processing for data structures and effective handling of nested loops, enabling it to tackle intricately structured programs. Second, ACInv delivers a fully automated, end-to-end solution for loop invariant generation, significantly reducing the need for manual intervention throughout the process. Lastly, while maintaining a competitive level of correctness, ACInv minimizes time overhead and optimizes the quality of the generated invariants.
To evaluate ACInv, we conducted a series of experiments. We conducted experiments on ACInv on datasets with and without data structures. The experiments showed that ACInv performed better than all current tools on datasets with data structures. On datasets without data structures, ACInv also performed close to the performance of the state-of-art tool AutoSpec [22]. Overall, ACInv can solve 21.08% more examples than AutoSpec.
@@
  • We have built a better automatic invariant generation analysis method based on LLM that can handle relatively complex data structures while maintaining a relatively high correctness rate with low local time consumption, and built a tool ACInv based on it.
  • We generated some predicate templates that may be useful for some commonly used data structures and other user-defined data structures on benchmarks such as  [23, 29] and SV-COMP [30, 31].
@@ 2.1 Loop Invariant
To provide a formal characterization of the invariant, we first abstract the loop and its invariants using Hoare logic [1] as follows:
$\frac{\{I\wedge B\}\ S\ \{I\}}{\{I\}\ \mathbf{while}\ (B)\ \mathbf{do}\ (S)\ \{I\wedge\neg B\}}$
Where we abstract the general loop structure in the form of pseudo-code into the expression $\mathbf{while}\ (B)\ \mathbf{do}\ (S)$, defining $B$ as the guard condition, $S$ as the loop body, and $I$ as the sought loop invariant. Given a loop invariant $I$, if the body $S$ preserves $I$ under the condition that both $I$ and $B$ hold, i.e., $\{I\wedge B\}\ S\ \{I\}$ is valid, then initiating the loop in a state satisfying $I$ ensures that upon termination, the state satisfies both the invariant $I$ and the negation of the guard condition $\neg B$, formally, $\{I\}\ \mathbf{while}\ (B)\ \mathbf{do}\ (S)\ \{I\wedge\neg B\}$. This rule underscores the critical role of identifying a suitable loop invariant $I$ for verifying loop behavior.
Figure 1: A simple C code example with loop invariants.
An example of loop invariants in ACSL The vast majority of invariant research is based on the C language, so this paper will also use C language code as the research object. In the following paper, we will use ANSI/ISO C Specification Language (ACSL) to describe the loop invariants of the C program, which can be automatically checked by the verification tool Frama-C [5] and its plugin WP. In ACSL, all the information needed is implemented as the annotation which begins with \\@ or is wrapped with \*@...*\, where loop invariants are annotated beginning with loop invariant.
Consider the example depicted in Fig 1. The provided C code snippet above demonstrates a simple while loop. The program starts by declaring an integer variable i and initializing it to 0. The while loop runs as long as the condition i < 30 holds true. During each iteration, the value of i is incremented by 1 (++i). When i reaches 30, the condition i < 30 becomes false, and the loop terminates. The C program implements a bounded iteration construct with a formally verified loop invariant $\mathtt{0\leq i\leq 30}$ specified in ACSL annotations, where the loop initializes $\mathtt{i}$ to 0 and increments it via pre-increment operation $\mathtt{++i}$ until the guard condition $\mathtt{i<30}$ becomes false. Through Hoare logic verification, the invariant satisfies initialization at entry ($\mathtt{i=0}\Rightarrow\mathtt{0\leq 0\leq 30}$), preservation under execution ($\mathtt{0\leq i\leq 30}\land\mathtt{i<30}\vdash\mathtt{0\leq i+1\leq 30}$ since $\mathtt{i\leq 29}$), and conclusion upon termination ($\mathtt{0\leq i\leq 30}\land\neg\mathtt{(i<30)}\equiv\mathtt{i=30}$).
What we still need to add and explain is, the invariant and predicate referenced herein, along with those mentioned subsequently, are incorporated into the original code in the form of annotations. However, these annotations differ significantly from conventional comments. Traditional comments are authored by programmers to articulate their design intentions informally. In contrast, predicates and invariants are employed to specify certain program properties rigorously; they serve to assist in program verification and similar tasks, and are expressed in a strict formal language based on the ACSL (ANSI/ISO C Specification Language). Consequently, they serve distinct purposes and are conventionally demarcated by specialized syntactic markers in practice.
@@ 2.2 Large Language Models (LLMs)
In recent years, the field of Natural Language Processing (NLP) has achieved significant advances. State-of-the-art models, particularly those built on the transformer architecture [32], have achieved remarkable success in addressing a wide range of natural language tasks. Large Language Models (LLMs), sophisticated neural networks trained on massive text corpora, have emerged as powerful tools for understanding and generating human-like language. By utilizing deep learning frameworks such as transformers, LLMs can effectively capture and model intricate linguistic patterns and relationships. These models represent a major leap forward in NLP, driven by the convergence of novel architectures, vast datasets, and enhanced computational resources. The development of large-scale pre-trained models like BERT (Bidirectional Encoder Representations from Transformers) [33] and GPT (Generative Pre-trained Transformer) [34] has led to substantial improvements across a variety of NLP tasks.
Prompting is a technique where specific input phrases or instructions, known as "prompts", are used to guide the LLM’s responses. Prompting leverages the model’s pre-existing knowledge, allowing for immediate, on-the-fly task execution without the need for additional training. This method is highly flexible and efficient, especially for exploratory or dynamic tasks, as it does not require any modifications to the model’s internal parameters. Previous work has demonstrated that techniques such as Chain-of-Thought (CoT) prompting [26] and few-shot learning [34] are highly effective in enhancing the performance of prompts. CoT prompting encourages the model to break down complex problems into intermediate reasoning steps, enabling more structured and logical responses. Few-shot learning involves providing the model with a small number of task-relevant examples within the prompt, which helps the model understand the specific format, context, or structure required for the task, even when it has not been explicitly fine-tuned for it. Drawing from various prompting techniques [25, 26, 27, 28], we designed a structured prompt framework to help us enhance computational efficiency in deploying LLMs for inference.
3 Methodology
3.1 Overview
@@
In Figure 2, we present an overview diagram illustrating our proposed approach. To address the aforementioned challenge, we introduce ACInv, an LLM-based tool for automatic loop invariant generation. ACInv’s process consists of three main phases: (1) Extraction Phase: ACInv initially extracts data structure and loop information from the target program. This involves performing basic control flow and value analyses, classifying loops, and gathering relevant variable information for each loop. (2) Generation Phase: During this phase, ACInv embeds the extracted data structure information into data structure prompts, querying LLMs to derive predicates that abstract key properties of the data structure. These predicates, along with loop-specific information, are then embedded into invariant prompts to generate loop invariants. If the current loop is not the first in the program, information from preceding loops, along with their generated invariants, will also be incorporated into the prompts. (3) Augmentation Phase: This section mainly consists of two components: the evaluator, and the optimizer, which are all composed of LLMs. The evaluator assesses the correctness of each generated invariant. For correct invariants, the optimizer attempts to strengthen them. For incorrect invariants, the optimizer weakens them to align with the correct program state. If an invariant cannot be refined to the correct level after $k$ iterations, it is discarded. In the remainder of this section, we will provide a detailed explanation of the entire methodology and its related intricacies.
3.2 Problem description